OpenStudy (anonymous):

Calculate the discriminant. x2 – x + 2 = 0 A. 9 B. –7 C. 7 D. –9

OpenStudy (anonymous):

Formula for the discriminant is b^2-4ac ax^2+bx+c is the form of the polynomial, thus a=1, b=-1, and c=2 (-1)^2-4(1)(2)= 1-8=-7 or B

OpenStudy (anonymous):

you there??

OpenStudy (anonymous):

It should be there? Just in case: Formula for the discriminant is b^2-4ac ax^2+bx+c is the form of the polynomial, thus a=1, b=-1, and c=2 (-1)^2-4(1)(2)= 1-8=-7 or B

OpenStudy (anonymous):

yea its there but which answer is it?

OpenStudy (anonymous):

The answer is B. -7

OpenStudy (anonymous):

thank you i have more. :P

OpenStudy (anonymous):

I can take a look (: